# Beyond Accuracy-Fairness: Stop evaluating bias mitigation methods solely on between-group metrics

### v1

@@ 4.2 Can we compare the bias mitigation methods?
Table 3: Results of the bias mitigation strategies on the five datasets. We report: the AUC score overall, and split over the protected group and the privileged group, the accuracy (ACC), statical parity difference (SPD), equal opportunity difference (EOD) and the positive rate (PR). Best values are highlighted in bold.
Dataset Metric ML model LFR DIR ADV MFC ROC EOP TO Adult AUC 0.843 0.623 0.85 0.847 0.826 0.843 0.843 0.843 $AUC^{pro},AUC^{pri}$ 0.811, 0.826 0.501, 0.508 0.824, 0.835 0.834, 0.843 0.789, 0.816 0.811, 0.826 0.811, 0.826 0.811, 0.826 ACC 0.806 0.766 0.825 0.82 0.81 0.728 0.77 0.793 SPD -0.26 -0.005 -0.164 -0.09 -0.139 -0.055 -0.046 -0.005 EOD -0.139 0.0 -0.181 -0.274 -0.278 -0.332 -0.283 -0.368 PR 0.247 0.005 0.174 0.161 0.178 0.393 0.151 0.167 Compas AUC 0.834 0.693 0.833 0.808 0.832 0.834 0.834 0.834 $AUC^{pro},AUC^{pri}$ 0.814, 0.821 0.588, 0.703 0.81, 0.823 0.809, 0.815 0.812, 0.821 0.814, 0.821 0.814, 0.821 0.814, 0.821 ACC 0.758 0.645 0.753 0.737 0.736 0.732 0.683 0.727 SPD -0.376 -0.866 -0.332 -0.064 -0.211 -0.031 -0.07 0.014 EOD -0.13 -0.055 -0.131 -0.231 -0.175 -0.233 -0.25 -0.262 PR 0.518 0.49 0.617 0.58 0.696 0.528 0.545 0.569 Dutch AUC 0.887 0.657 0.887 0.874 0.883 0.887 0.887 0.887 $AUC^{pro},AUC^{pri}$ 0.884, 0.848 0.499, 0.498 0.884, 0.849 0.881, 0.847 0.89, 0.852 0.884, 0.848 0.884, 0.848 0.884, 0.848 ACC 0.812 0.476 0.786 0.768 0.695 0.776 0.754 0.763 SPD -0.318 0.0 -0.432 -0.171 -0.394 -0.066 -0.159 -0.02 EOD -0.026 -0.315 -0.047 -0.073 -0.024 -0.217 -0.243 -0.254 PR 0.416 1.0 0.586 0.303 0.203 0.395 0.45 0.396 Law AUC 0.882 0.83 0.883 0.879 0.122 0.882 0.882 0.882 $AUC^{pro},AUC^{pri}$ 0.848, 0.864 0.803, 0.792 0.853, 0.864 0.843, 0.862 0.146, 0.142 0.848, 0.864 0.848, 0.864 0.848, 0.864 ACC 0.903 0.897 0.903 0.901 0.22 0.772 0.879 0.892 SPD -0.197 -0.207 -0.184 -0.141 0.494 -0.044 -0.021 -0.002 EOD -0.111 -0.128 -0.122 -0.143 -0.151 -0.104 -0.197 -0.21 PR 0.954 0.967 0.961 0.965 0.289 0.711 0.961 0.979 Student AUC 0.803 0.693 0.817 0.772 0.762 0.803 0.803 0.803 $AUC^{pro},AUC^{pri}$ 0.819, 0.788 0.689, 0.642 0.819, 0.814 0.757, 0.779 0.781, 0.794 0.819, 0.788 0.819, 0.788 0.819, 0.788 ACC 0.759 0.626 0.759 0.703 0.738 0.697 0.728 0.728 SPD -0.104 -0.72 -0.065 -0.573 0.031 0.076 -0.016 0.071 EOD -0.166 -0.063 -0.202 0.027 -0.283 -0.191 -0.205 -0.246 PR 0.585 0.697 0.574 0.662 0.677 0.492 0.595 0.605
We present the results of performance and fairness metrics for bias mitigation methods across five datasets in Table 3.
@@
Aligning with existing literature, we observe that fairness metrics such as SPD and EOD often yield conflicting results (Kleinberg et al., 2016). It is rarely the same method that returns the best results for both metrics. Furthermore, no single method consistently outperforms all others for one of the metrics. However, DIR stands out for its excellent AUC performance, aligning with its design to preserve rank-ordering within groups (Feldman et al., 2015).
@@
When comparing bias mitigation methods, conventional benchmarking studies often focus on accuracy and fairness (Chen et al., 2023; Zemel et al., 2013). However, relying solely on accuracy, which uses prediction labels at a specific threshold, may lead to inappropriate comparisons. There are two reasons why using the prediction labels is not a suitable approach: First of all, the fairness metrics are not yet satisfied, and secondly, different mitigation methods can result in varying positive rates. Both reasons would lead to an additional altering of the prediction labels post-hoc, so comparing them at this stage does not seem sensible. A more comprehensive approach involves assessing the performance of the prediction scores.
@@
Regarding the second concern, Table 3 reveals significant positive rate disparities among different strategies. This inconsistency poses challenges in real-world applications, where a fixed or reasonably bounded positive rate is typically expected. For example, when attempting to satisfy one of the fairness metrics, practitioners might consider LFR in the Adult Income dataset, despite a slight accuracy loss. However, this choice results in an unexpectedly low positive rate (0.5%), deviating significantly from other methods. Similarly, for the Dutch dataset, MFC leads to the best value for EOD, but it only has a positive rate of 20.3%, while the initial model has a positive rate of 41.6%. DIR leads to the best AUC, but has a positive rate of 58.6%. We argue against treating methods with significantly different positive rates as comparable situations. In practice, most real-world applications will have a relatively fixed positive rate, and bias mitigation methods must be adapted accordingly.
@@
Both these concerns can effectively be addressed by adjusting the classification threshold(s) of the predictive model. Consequently, it is more logical to evaluate these mitigation methods based on their prediction scores, rather than on their prediction labels, which are still subject to change. We advocate to compare these mitigation methods based on the AUC score, to assess how well the indiduals are ranked within each group, and to generate the prediction labels post-hoc, based on the chosen fairness metric and practical constraints.
@@
It is crucial not only to examine the overall AUC score across the entire population, but also to scrutinize the AUC scores disaggregated by different groups. Certain bias mitigation methods may yield the optimal ranking for one group but not for another. This prompts the need for decision-making: should preference be given to the overall best ranking or to narrowing the gap in rankings between the protected and privileged groups? Alternatively, deploying two separate models could be considered to ensure the best ranking for each subgroup. Given these results, one should evaluate whether a slight improvement in AUC justifies the adoption of distinct models. Additionally, note that for both the Adult and the Dutch datasets, the LFR method reduces the within-group ranking to an essentially random ordering. In any scenario, deploying a bias mitigation method that decrease the AUC score for every subgroup is counterproductive when we want to optimize for both performance and fairness. [^5] Note that we operate under the assumption that there is no label bias. Unfortunately, this is a common outcome in practice, as many bias mitigation methods may compromise AUC scores across subgroups in an attempt to adhere to a fairness metric (Mittelstadt et al., 2023).
@@
Note that the goal of this comparison is not to declare the superior performance of one of the mitigation methods. For this, a more comprehensive benchmark study is needed with a larger number of datasets, machine learning models, and extensive tuning of each bias mitigation method. This was already the goal of multiple other benchmark studies (Chen et al., 2023; Hort et al., 2021; Reddy, 2022). Our primary goal was to emphasize the inadequacy of the current way of benchmarking these methods with each other, based on prediction labels that should still be subject to change.
4.3 How should bias mitigation methods be implemented?
Figure 3: Proposed framework to mitigate bias in machine learning models. Note that all actions mentioned under Step 1 are optional, and can be used in combination with each other.
We propose a novel framework for implementing bias mitigation strategies, outlined in Figure 3. In the initial step, the focus should be on optimizing the ranking accuracy for each subgroup independently. Evaluating the AUC score for each subgroup is crucial, as a lower AUC score for one subgroup can arise from various factors. For example, there may be insufficient data available for this subgroup, limiting the model’s ability to accurately capture its characteristics. Additionally, the data quality for the protected group could be notably lower, introducing noise or inaccuracies into the learning process. Moreover, the model might learn patterns based on historical inequalities that are currently undergoing changes. The evolving nature of societal dynamics can lead to shifts in behavior and characteristics within subgroups, impacting the model’s predictive accuracy. Another potential reason is that the model learned patterns primarily from the privileged group, often the majority group, which do not accurately resemble the behavior of the protected group. Whatever the cause, the first step should be to enhance the ranking accuracy within each group. This can be done by selecting more data from that subgroup, improving data quality, employing feature engineering, deploying separate models for each group, removing sensitive attributes, or utilizing preprocessing or inprocessing mitigation techniques. Note that all these proposed interventions are optional, as they can also lead to a lower AUC score. This list is by no means exhaustive, and a combination of multiple strategies may be the most effective.
Subsequently, the second step involves selecting the top-ranked individuals from each group, aligning with the desired fairness metrics and the applicable industry constraints. For instance, in the context of college admissions, adhering to demographic parity and having the capacity to admit approximately 10% of applicants could entail selecting the 10% highest-ranked instances from each subgroup. This two-step framework aims to not only enhance the accuracy of rankings within subgroups but also align the selection process with predefined fairness criteria and practical considerations.
Figure 4: Accuracy when selecting the top ranked individuals using the prediction scores produced by MFC and the initial ML model. This analysis is done for for different positive rates.
Let’s walk through the application of our framework using the Dutch dataset. Table 3 reveals that the best ranking results for each subgroup are achieved with the MFC method. However, this improvement comes at the cost of a significant decrease in accuracy, dropping from 0.812 to 0.695. It’s important to note that comparing these situations directly may not be entirely fair, given that the positive rate of the initial machine learning model is more than double than the positive rate after applying MFC. Additionally, in both cases, the fairness metrics are not satisfied.
Now, let’s apply our framework and consider enforcing demographic parity. We select the top-ranked individuals from each group based on the prediction scores after applying MFC, and for the sake of comparison also after applying the initial machine learning model. We illustrate the resulting accuracy for various positive rates in Figure 4. Notably, for all tested positive rates, the accuracy achieved using the scores produced by MFC is as high as or slightly higher than the scores from the initial machine learning model.
Two key points should be highlighted in interpreting these results. First, we refrain from claiming that the ranking produced by MFC is optimal. Achieving optimality would require testing more machine learning models, finetuning all ML models and mitigation methods, or testing some of the other suggestions made in Figure 3. Our aim is to underscore that relying solely on accuracy metrics can lead to misinterpretations, as our subsequent steps should use the prediction scores rather than the prediction labels.
Additionally, it’s crucial to acknowledge that in this specific case, we selected demographic parity as the fairness metric, which is easily enforced by selecting the top-ranked individuals from each subgroup. Other fairness metrics may require using postprocessing mitigation methods or involve solving an optimization problem that considers both the fairness metric and the desired positive rate.
@@ 5 Discussion
In this paper, our initial aim is to provide clarity on the specific impact of bias mitigation methods on various instances. A crucial aspect involves comparing the ranking of instances before and after the application of bias mitigation methods. We observe that preprocessing and inprocessing methods introduce changes to the ranking, whereas postprocessing methods maintain the original ranking and simply adjust labels based on it.
Furthermore, we contend that the current approach to comparing bias mitigation methods is insufficient due to the sole focus on between-group metrics and the comparison of disparate scenarios. To address this limitation, we propose a novel evaluation framework. The framework comprises two key steps: Firstly, establish the most accurate ranking for each subgroup. Secondly, select the highest-ranked individuals from each subgroup, guided by fairness metrics and industry-specific constraints. This refined framework aims to provide a more nuanced and comprehensive assessment of bias mitigation methods, aligning with the intricacies of real-world applications. This aligns with a paradigm shift from between-group fairness to within-group fairness, where the goal is to do as well as possible for each subgroup. By incorporating the existing capacity in positive instances as a constraint within our framework, we effectively prevent scenarios where the positive rate is excessively high and unrealistic, as well as situations where capacity remains underutilized. Regrettably, this underutilization is a frequent outcome in the application of bias mitigation strategies, where the positive rate of the privileged group is often reduced to match that of the underprivileged group, resulting in no significant improvement for either group (Mittelstadt et al., 2023).
Although we posit that for optimal fairness and performance, pre- and inprocessing methods should only be employed if they enhance subgroup rankings, there are scenarios where employing these methods might be beneficial, even if they result in less favorable subgroup rankings. Our framework can lead to situations where two otherwise identical entities are treated differently based solely on a sensitive attribute, a practice that may be unlawful in certain contexts. Furthermore, these methods can be advantageous in situations where applying distinct thresholds for different groups is impractical, or in cases where the decision-making body does not have access to the sensitive attribute. Such circumstances warrant the consideration of these methods despite their potential drawbacks in subgroup ranking. However, even in these cases, it is imperative to evaluate them at realistic positive rates.
@@
An obvious limitation of our framework is that it operates under the assumption of having access to a static sensitive attribute, and it inherently faces difficulties in scenarios where these assumptions may not hold. For instance, the assumption of the sensitive attribute being static overlooks the evolving nature of certain attributes, such as gender, where people can identify as other categories over time. Another limitation arises when considering the assumption of unrestricted access to the sensitive attribute. In reality, legal and ethical considerations may impose constraints on obtaining or utilizing certain sensitive information (Haeri and Zweig, 2020; Veale and Binns, 2017; Johnson, 2021; Holstein et al., 2019). For instance, privacy laws and regulations may restrict the collection or use of specific attributes, further complicating the application of this framework in practice. These limitations underscore the need for ongoing refinement and adaptation of fairness frameworks to accommodate the complexities and nuances inherent in real-world scenarios.
Acknowledgments
This research was funded by Flemish Research Foundation (Grant number 11N7723N).
Appendix A Appendices
A.1 Score distributions for the other datasets

### v3

@@ 5.2 Can we compare the bias mitigation methods?
Table 3: Results of the bias mitigation strategies on the five datasets. We report: the AUC score overall, and split over the protected group and the privileged group, the accuracy (ACC), statical parity difference (SPD), equal opportunity difference (EOD) and the positive decision rate (PDR). Best values are highlighted in bold.
Dataset Metric ML model LFR DIR ADV MFC ROC EOP TO Adult AUC 0.843 0.623 0.85 0.847 0.826 0.843 0.843 0.843 $AUC^{pro},AUC^{pri}$ 0.811, 0.826 0.501, 0.508 0.824, 0.835 0.834, 0.843 0.789, 0.816 0.811, 0.826 0.811, 0.826 0.811, 0.826 ACC 0.806 0.766 0.825 0.82 0.81 0.728 0.77 0.793 SPD -0.26 -0.005 -0.164 -0.09 -0.139 -0.055 -0.046 -0.005 EOD -0.139 0.0 -0.181 -0.274 -0.278 -0.332 -0.283 -0.368 PDR 0.247 0.005 0.174 0.161 0.178 0.393 0.151 0.167 Compas AUC 0.834 0.693 0.833 0.808 0.832 0.834 0.834 0.834 $AUC^{pro},AUC^{pri}$ 0.814, 0.821 0.588, 0.703 0.81, 0.823 0.809, 0.815 0.812, 0.821 0.814, 0.821 0.814, 0.821 0.814, 0.821 ACC 0.758 0.645 0.753 0.737 0.736 0.732 0.683 0.727 SPD -0.376 -0.866 -0.332 -0.064 -0.211 -0.031 -0.07 0.014 EOD -0.13 -0.055 -0.131 -0.231 -0.175 -0.233 -0.25 -0.262 PDR 0.518 0.49 0.617 0.58 0.696 0.528 0.545 0.569 Dutch AUC 0.887 0.657 0.887 0.874 0.883 0.887 0.887 0.887 $AUC^{pro},AUC^{pri}$ 0.884, 0.848 0.499, 0.498 0.884, 0.849 0.881, 0.847 0.89, 0.852 0.884, 0.848 0.884, 0.848 0.884, 0.848 ACC 0.812 0.476 0.786 0.768 0.695 0.776 0.754 0.763 SPD -0.318 0.0 -0.432 -0.171 -0.394 -0.066 -0.159 -0.02 EOD -0.026 -0.315 -0.047 -0.073 -0.024 -0.217 -0.243 -0.254 PDR 0.416 1.0 0.586 0.303 0.203 0.395 0.45 0.396 Law AUC 0.882 0.83 0.883 0.879 0.122 0.882 0.882 0.882 $AUC^{pro},AUC^{pri}$ 0.848, 0.864 0.803, 0.792 0.853, 0.864 0.843, 0.862 0.146, 0.142 0.848, 0.864 0.848, 0.864 0.848, 0.864 ACC 0.903 0.897 0.903 0.901 0.22 0.772 0.879 0.892 SPD -0.197 -0.207 -0.184 -0.141 0.494 -0.044 -0.021 -0.002 EOD -0.111 -0.128 -0.122 -0.143 -0.151 -0.104 -0.197 -0.21 PDR 0.954 0.967 0.961 0.965 0.289 0.711 0.961 0.979 Student AUC 0.803 0.693 0.817 0.772 0.762 0.803 0.803 0.803 $AUC^{pro},AUC^{pri}$ 0.819, 0.788 0.689, 0.642 0.819, 0.814 0.757, 0.779 0.781, 0.794 0.819, 0.788 0.819, 0.788 0.819, 0.788 ACC 0.759 0.626 0.759 0.703 0.738 0.697 0.728 0.728 SPD -0.104 -0.72 -0.065 -0.573 0.031 0.076 -0.016 0.071 EOD -0.166 -0.063 -0.202 0.027 -0.283 -0.191 -0.205 -0.246 PDR 0.585 0.697 0.574 0.662 0.677 0.492 0.595 0.605
@@
We present the results of the performance and fairness metrics for all bias mitigation methods across five datasets in Table 3. Aligning with existing literature, we observe that fairness metrics such as SPD and EOD often yield conflicting results (Kleinberg et al., 2016). It is rarely the same method that returns the best results for both metrics. Furthermore, no single method consistently outperforms all others for one of the metrics. However, DIR stands out for its excellent AUC performance, aligning with its design to preserve rank-ordering within groups (Feldman et al., 2015).
Can we use the prediction labels to compare the methods?
@@
When comparing bias mitigation methods, conventional benchmarking studies often focus on accuracy and fairness (Chen et al., 2023; Zemel et al., 2013). However, relying solely on accuracy, which uses prediction labels at a specific threshold, may lead to inappropriate comparisons. As mentioned, there are two reasons why using the prediction labels is not a suitable approach: First of all, the fairness metrics are not yet satisfied, and secondly, different mitigation methods can result in varying positive rates. Both reasons would lead to an additional altering of the prediction labels post-hoc, so comparing them at this stage does not seem sensible. A more comprehensive approach involves assessing the performance of the prediction scores, or comparing the prediction labels when the thresholds have been modified to address both industry constraints and fairness considerations.
@@
Regarding the second concern, Table 3 reveals significant positive decision rate disparities among different strategies. This inconsistency poses challenges in real-world applications, where a fixed or reasonably bounded positive decision rate is typically expected (Kwegyir-Aggrey et al., 2023). For example, when attempting to satisfy one of the fairness metrics, practitioners might consider LFR in the Adult Income dataset, despite a slight accuracy loss. However, this choice results in an unexpectedly low positive decision rate (0.5%), deviating significantly from other methods. This method just predicts almost every instance as negative, which results in a satisfaction of the fairness metrics, but is not realistic in the real-word. Similarly, for the Dutch dataset, MFC leads to the best value for EOD, but it only has a positive decision rate of 20.3%, while the initial model has a positive decision rate of 41.6%. DIR leads to the best AUC, but has a positive decision rate of 58.6%. This demonstrates the second side effect of using bias mitigation methods, namely an altering of the global selection rate. We argue against treating methods with significantly different positive decision rates as comparable situations. In practice, most real-world applications will have a relatively fixed positive decision rate, and bias mitigation methods must be adapted accordingly.
Evaluate by using the prediction scores
@@
Both these concerns can effectively be addressed by adjusting the classification threshold(s) of the predictive model. Consequently, it makes more sense to evaluate these mitigation methods based on their prediction scores, rather than on their prediction labels, which are still subject to change. We advocate to compare these mitigation methods based on the AUC score, to assess how well the individuals are ranked within each group, and to generate the prediction labels post-hoc, based on the chosen fairness metric and practical constraints.
@@
It is important not only to examine the overall AUC score across the entire population, but also the AUC scores disaggregated by different groups. Certain bias mitigation methods may yield the optimal ranking for one group but not for another. This prompts the need for decision-making: should preference be given to the overall best ranking or to narrowing the gap in rankings between the protected and privileged groups? Alternatively, deploying two separate models could be considered to ensure the best ranking for each subgroup. Given these results, one should evaluate whether a slight improvement in AUC justifies the adoption of distinct models. Additionally, note that for both the Adult and the Dutch datasets, the LFR method reduces the within-group ranking to an essentially random ordering. In any scenario, deploying a bias mitigation method that decreases the AUC score for every subgroup seems counterproductive when we want to optimize for both performance and fairness. [^13] Note that we operate under the assumption that there is no label bias. Unfortunately, this is a common outcome in practice, as many bias mitigation methods may compromise AUC scores across subgroups in an attempt to adhere to a fairness metric (Mittelstadt et al., 2023).
@@
Note that the goal of this comparison is not to declare the superior performance of one of the mitigation methods. For this, a more comprehensive benchmark study is needed with a larger number of datasets, machine learning models, and extensive tuning of each bias mitigation method. This was already the goal of multiple other benchmark studies (Chen et al., 2023; Hort et al., 2021; Reddy, 2022). Our primary goal was to show some of the unintended side effects of bias mitigation methods, such as an altering of the global selection rate and the effects within each group, and to emphasize the inadequacy of the current way of benchmarking these methods with each other, based on prediction labels that are still be subject to change.
@@ 6 Discussion
In this study , we demonstrate two side effects of deploying bias mitigation methods that are not (sufficiently) taken into account. First, bias mitigation methods not only introduce significant changes between-groups but also within-groups and these changes currently go unnoticed. This reranking process is not necessarily an issue, but deserves more attention. Second, they lead to an altering of the global selection rate which results in very different scenarios. We argue that the current approach to comparing bias mitigation methods is insufficient due to the sole focus on the output of the prediction model without taking into account the decision-making system (Scantamburlo et al., 2024).
Regarding the first point, is this reranking process necessarily a bad thing? We notice in Table 3 that the bias mitigation methods can result in slightly better, worse or approximately the same ranking accuracy (measured by the AUC) as the initial machine learning model. If the ranking improves after using the bias mitigation method, there is no issue, as the method will lead to better rankings that should also be more fair. But what if the ranking accuracy is significantly worse than the the output of the initial machine learning model? Is this always undesirable? Not necessarily. Until, and consistent with most literature in fair machine learning, we assumed that we only want to remove bias between groups. However, if we assume there is also bias within each group (so the Affirmative Action Theorem does not hold) and this is undesirable, then it makes sense to use mitigation methods that also shift the ranking within each group, even if they appear worse with respect to the target label.
If we assume that there is no within-group bias, we propose that for the dual objectives of fairness and accuracy optimization, preprocessing and inprocessing mitigation methods should be adopted only if they improve subgroup rankings. Should they fail to do so, we posit that utilizing the ranking produced by the original machine learning model and then applying Threshold Optimization post-hoc is sufficient and optimal, as it directly addresses the fairness measures without unnecessary within-group reranking. However, this practice is not always possible, as this can lead to situations where two otherwise identical entities are treated differently based solely on a sensitive attribute, a practice that may be unlawful in certain contexts. Furthermore, we can not always assume that the decision-making body has access to the sensitive attribute, and thus is able to do these fairness interventions post-hoc. In those settings, it might be better to use preprocessing or inprocessing bias mitigation methods, even if they worsen this trade-off.
Finally, what if the performance of the ranking stays approximately the same, but the ranking itself will be significantly different from the ranking of the initial machine learning model (as measured in Table 2)? Is this arbitrariness a problem? The opinions on this differ. The literature on ‘predictive multiplicity’ (Marx et al., 2020) or ‘model multiplicity’ (Black et al., 2022) discusses the situation where there exist many possible models with similar predictive performance but slightly different decisions on individuals, which is comparable to our situation. They argue that multiplicity should be reduced by removing some of the variance that leads to diverging predictions (Cooper et al., 2023). As we can see, there is no easy answer to this question, and we look forward to more debate about this topic.
Limitations
A first limitation of our study is the assumption of no label bias. This leads us to quite a straddle, which is common in fair machine learning. On the one hand, we presume that our labels are correct to ensure the reliability of our metrics. On the other hand, we recognize that models might need adjustments to correct for bias. This make sense for equal opportunity, where the goal is to ensure that all groups have equal true positive rates, thus requiring a classifier that does not make the situation worse. For demographic parity, the issue is more complex: why would we want to deviate from the labels if they are correct? Nonetheless, adjustments may be necessary to align with external policy or legal requirements.
@@
A second limitation of our study is that it operates under the assumption of having access to a static sensitive attribute, and it will face difficulties in scenarios where these assumptions may not hold. For instance, the assumption of the sensitive attribute being static overlooks the evolving nature of certain attributes, such as gender, where people can identify as other categories over time. Another limitation arises when considering the assumption of unrestricted access to the sensitive attribute. In reality, legal and ethical considerations may impose constraints on obtaining or utilizing certain sensitive information (Haeri and Zweig, 2020; Veale and Binns, 2017; Johnson, 2021; Holstein et al., 2019). For instance, privacy laws and regulations may restrict the collection or use of specific attributes, further complicating research in fairness.
Appendix A Proof of Theorem 3.7
Theorem A.1.
The Affirmative Action assumption holds if and only if for all $a\in A$ it holds that for almost all $(x,a),(x^{\prime},a)\in X\times\{a\}$
$p(Y=1|X=x,A=a)\leq p(Y=1|X=x^{\prime},A=a)\Rightarrow S(x,a)\leq S(x^{\prime},a)$
Proof.
()($\Rightarrow$)( ⇒ ) Let’s suppose by way of contradiction that the implication is not true for some sensitive attribute $a\in A$. Formally this means that there exists a non-zero subset $U\times\{a\}\times U^{\prime}\times\{a\}\subseteq X\times\{a\}\times X\times\{a\}$ such that for all $x\in U,x^{\prime}\in U^{\prime}$ we have
$p(Y=1|X=x,A=a)\leq p(Y=1|X=x^{\prime},A=a)\quad\text{ and}\quad S(x,a)>S(x^{\prime},a)$
Let’s now take a threshold $t$ such that
$\sup_{x\in U}p(Y=1|X=x,A=a)\leq t\leq\inf_{x^{\prime}\in U^{\prime}}p(Y=1|X=x^{\prime},A=a)$
such threshold $t$ belongs in the interval $(0,1)$ since $p$ is supposed continuous and since $U\times\{a\}$ and $U^{\prime}\times\{a\}$ have non-zero measure. Let $\hat{Y}$ be a decision associated with the threshold $t$, that is
$\hat{Y}(x,a)=\begin{cases}1&\text{if }p(Y=1|X=x,A=a)>t\\ 0&\text{otherwise}\end{cases}$
In particular, almost everywhere we have $\hat{Y}(U\times\{a\})=0$ and $\hat{Y}(U^{\prime}\times\{a\})=1$. We claim that such decision cannot be optimal according to $S(x,a)$: if it were, it would exist a threshold $t^{\prime}$ such that almost everywhere
$\hat{Y}(x,a)=\begin{cases}1&\text{if }S(x,a)>t^{\prime}\\ 0&\text{otherwise}\end{cases}$
Then for almost all $x\in U$ and $x^{\prime}\in U^{\prime}$
$t^{\prime}\leq S(x^{\prime},a)<S(x,a)\leq t^{\prime}$
which is a contradiction.
()($\Leftarrow$)( ⇐ ) From the hypothesis almost everywhere we have
$p(Y=1|X=x,A=a)=p(Y=1|X=x^{\prime},A=a)\Rightarrow S(x,a)=S(x^{\prime},a)$
So people from the same sensitive group with the same fair probabilities are assigned the same score as well. So we can define a bias function $\beta$ such that almost everywhere
$\beta\left(p(Y=1|X=x,A=a),a\right)=S(x,a)$
The function $\beta$ takes as input the fair probability and the sensitive attribute and returns the score. By hypothesis, $\beta$ is a non-decreasing function of the fair probability. If we now take a maximal decision $\hat{Y}$ according to $p$, then for all $a\in A$ we can find a decision $\hat{Y}_{a}\colon X\times\{a\}\to\{0,1\}$ defined as follows
$\hat{Y}_{a}(x,a)=\begin{cases}1&\text{if }S(x,a)>\beta\left(t,a\right)\\ 0&\text{otherwise}\end{cases}$
which are maximal decisions according to $S(x,a)$ and such that $\hat{Y}=\hat{Y}_{0}\cup\hat{Y}_{1}$. ∎
Appendix B Correlation plots
Figure 2: Correlation plots
Appendix C Score distributions for the other datasets